theoretical maximum percent recovery

The solubility of benzoic water in acid at 0 degrees C is .90gm/100ml. In a recrystallization of benzoic acid, 22.5gm benzoic acid is dissolved in 450ml boiling water.This solution is allowed to cool to room temperature and then placed in an ice bath. What is the theoretical maximum percent recovery?
Answer #1

At 0 degrees C:

100 mL of water contains 0.90 g of benzoic acid

450 mL of water contains 450/100 x 0.90 = 4.05 g of benzoic acid

Mass of benzoic acid recovered = 22.5 - 4.05 = 18.45 g

Theoretical maximum percent recovery

= mass recovered/original mass dissolved x 100%

= 18.45/22.5 x 100%

= 82%

answered by: Tati